perimeter of a rectangle is 204 inches. the length exceeds the width by 62 inches. what is the length and width?
Answer by mananth(16949)

width be x
length = x+62
..
2*(L+W)=204
2*(x+x+62)=204
2*(2x+62)=204
divide by2
2x+62=102
2x=102-62
2x=40
/2
x= 20 the width
length = x+62 = 20+62
=82
...
.
m.ananth@hotmail.ca